OpenStudy (anonymous):

5sqrt x^10 simplify using rational exponents.

OpenStudy (anonymous):

sqrt(something) = (something)^(1/2) so u can write 5 (x^(1/2))^10 which is easy...

OpenStudy (anonymous):

\[\sqrt[5]{x ^{10}?}\]

OpenStudy (anonymous):

This is how the problem should look minus the ?

OpenStudy (anonymous):

Oh the 5th root, well how was I supposed to know that...? Anyway, same thing: nthrt(something) = (something)^(1/n) so u can write (x^(1/5))^10 which is easy...

OpenStudy (anonymous):

ok hang on would it be (x^10)^1/5 ?

OpenStudy (anonymous):

Powers, makes no difference which way round u put them, you are multiplying.

OpenStudy (anonymous):

\[(x ^{10})^{1/5}\]

OpenStudy (anonymous):

The fifth root of x to the 10th is the same as the tenth root of x to the fifth.

OpenStudy (anonymous):

ok I am really sorry , but is my answer right? I am confused

OpenStudy (anonymous):

To simplify powers like (a^b)^c you multiply the exponents so a^(bc). In this case, x^(10*1/5) = x^2

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

Oh ok, Thank you so much!!!

OpenStudy (anonymous):

U r welcome.